### v1

@@ 1 Introduction
There is a long tradition in labor economics of seeing jobs as more than pay (e.g., Brown, 1980; Duncan and Holmlund, 1983; Rosen, 1986). Expressions commonly used in the literature such as “compensating differentials,” “amenities,” or “workplace differentiation” all conceptualize the notion that jobs can be represented as a multidimensional bundle of attributes that workers trade off when considering opportunities with alternative employers. Such trade-offs can be substantial. Recent experimental evidence shows that workers at a US call center are willing to give up 8 percent of their salary for the option to work from home (Mas and Pallais, 2017), while undergraduates at a top US university would accept a 5.1 percent salary cut for a job offering the possibility to work part-time (Wiswall and Zafar, 2018).
What is unclear from this evidence, however, is how workers gather information on job attributes—both pay and non-pay—when exploring their options in the labor market. One important search channel through which workers can learn about these attributes is the content of job postings. [^4] Recently, Carrillo-Tudela et al. (2023) provide survey evidence from Germany showing that 55.3 percent of firms used job postings in their latest hire, while 88.1 percent of workers used job postings in their search process. Notably, however, job postings are not exclusive of other search methods, such as social networks and recruitment agencies. Carrillo-Tudela et al. (2023) find that firms and workers use, respectively, 1.9 and 2.3 channels on average. As employers ultimately decide on the content that they wish to mention in publicly advertised job postings, it remains an open question whether such content in fact reflects the actual pay and amenities associated with different employers.
@@
To characterize the distribution of job attributes advertised by employers in job ads, we use the near-universe of publicly posted vacancies in Norway between 2002 and 2019. A key contribution of our paper is to systematically extract job attributes from these texts. Using tools from natural language processing, we identify close to fifty commonly advertised pay and non-pay attributes, such as “competitive pay,” “flexible work hours,” and “nice work environment.” We build a large collection of unique expressions associated with each attribute, which we make publicly available online for other researchers to use. [^5] The full list of expressions is available in an online repository, accessible through this link. This approach allows us to identify a large number of job attributes and to transparently map each attribute to many alternative expressions.
@@
A key strength of our data set is that each ad can be directly traced to the establishment posting the job. To shed light on how different employers advertise different types of information in publicly posted job ads and to assess the information content of these publicly advertised job attributes, we estimate workers’ valuation of employers using job flows and wages observed in matched employer-employee data. Building on prior work by Sorkin (2018) and Morchio and Moser (2024), we use a structural revealed preference approach and obtain measures of workers’ valuation of different employers. Within our framework, this valuation can be further decomposed into a pay component, a non-pay component, and a job security component. Measuring these components separately matters because many of the attributes advertised in job listings, such as workplace quality, are not necessarily reflected in pay.
@@
We show that the content of job ads has predictive power for workers’ valuation of employers. Notably, we find that the attributes we detect in job ads can explain about 60 percent of the variation in overall employer values and employer pay premiums, and about 50 percent of non-pay and 40 percent of job security differences across employers. Interestingly, job ads have predictive power over and above commonly observed job characteristics. Once we remove variation in employer values that can be attributed to flexible industry, occupation, and location controls, job attributes listed in ads can still explain up to 20 percent of the remaining variation across employers. In this regard, we find that attributes related to having flexible working hours and shift work are especially important predictors. Further, controlling for industry, occupation and location, we show that employers with higher posted pay premiums and amenity values in job ads have shorter average duration of posted vacancies, consistent with the notion that such employers are more attractive.
Finally, we use our model estimates to gauge the impact on worker mobility of having access to the content of job ads. We predict the counterfactual value of employers under alternative assumptions on workers’ information sets and derive the implied mobility flows in each case. In our baseline, we find that aggregate job-to-job transitions to more highly-valued employers increase by 1.6 percent in an economy where workers have access to information on job attributes in job ads, as compared to a benchmark economy where only “one-liner” vacancies listing occupation, industry, and location are posted. This increase corresponds to 25 percent of the additional mobility towards better employers in the full-information model. Our results further suggest that the information in job ads can contribute to men and women self-selecting into distinct employers based on their preferences.
@@
Our results also have implications for the design of policies related to the information contained in job ads. There is a current policy drive to improve pay transparency within organisations in many countries, and several jurisdictions are also implementing policies to promote pay transparency in job openings Cullen (2024). [^6] As an example, in November 2022 New York City mandated employers to systematically include a salary bracket in their job ads. Taken together, our results on the prevalence of specific job attributes in vacancy texts and on the type of attributes advertised by better employers can be helpful to inform further policies in that direction.
@@
We contribute to this literature by instead considering a broad set of pay and non-pay attributes that are advertised by employers in job listings. In a similar vein, several recent contributions focus on a single specific attribute, such as the type of flexible work arrangements advertised in job ads (Adams-Prassl et al., 2020, 2023a), the possibility to work remotely (Hansen et al., 2023), and whether the employer offers training on the job (Adams-Prassl et al., 2023b). We add to these important contributions by retrieving a comprehensive set of job attributes advertised in the text of job openings, which encompasses flexible work arrangements, remote work and on-the-job training, but also what employers communicate about compensation or the quality of the workplace in their vacancies. In a similar spirit, Sockin (2022) uses text analysis to retrieve information on a large set of amenities from employer reviews posted by workers, rather than using vacancy texts. We see our results using vacancy texts as complementary since they instead reflect the information advertised by employers. [^7] In other recent and related work, Arold et al. (2024) and Lagos (2024) use text analysis methods to retrieve information on amenities contained in collective bargaining agreement texts.
@@
A large body of theoretical and empirical work has emphasized that non-pay attributes are important determinants of how much workers value alternative employers. One strand of this literature elicits workers’ preferences over specific attributes in survey experiments or quasi-experimental settings. Examples of non-pay attributes analyzed using this approach include workplace flexibility (Mas and Pallais, 2017; Wiswall and Zafar, 2018; Drake et al., 2023), job security (Datta, 2019), commuting time (Le Barbanchon et al., 2021), working from home (Nagler et al., 2022; Lewandowski et al., 2022), and shift work (Desiere and Walter, 2023), among others. Our paper instead provides a composite measure of employer amenity value that is linked to a large number of non-pay job attributes from job ads.
@@
We follow the second strand of this literature, which recovers structural estimates of the non-pay value of employers by modelling where workers choose to work in a labor market with heterogeneous employers. In this approach, estimates of the non-pay value of employers are obtained by reconciling workers’ choice of employer with their level of pay (Rosen, 1986). The choice of workers are captured using either the size of employers in a static setting (e.g., Card et al., 2018; Lamadon et al., 2022) or mobility flows between employers (e.g., Sorkin, 2018; Taber and Vejlin, 2020; Morchio and Moser, 2024). We rely on the latter approach since it can be recast in a job search framework, allowing us to further disentangle the amenity value of employers from differences in their recruitment effort. Relative to these important studies, we link the estimates of workers’ valuation of employers to the attributes advertised in job openings and quantify the value of this information. [^8] Morchio and Moser (2024) also relate their estimates of the amenity values of employers to some of the workplace characteristics observed in their data, such as “workplace hazards” and “working hours flexibility”.
@@
We also contribute to the expanding body of work that aims at unpacking search frictions in the labor market. Jäger et al. (forthcoming) elicit the beliefs of workers on their reemployment wage if they were to lose their job and compare these wages to actual labor market outcomes. Horton et al. (2024) and Belot et al. (2022) design experiments within a controlled environment to study how job seekers respond to variation in the wage featured in job vacancies. Our approach adds to these studies in that we focus on the texts of actual job postings as a source of information on alternative employers. Our evidence also relates to recent studies on how wages relate to the duration of posted vacancies (e.g., Mueller et al., 2024; Bassier et al., 2023; Faberman and Menzio, 2018). [^9] Several recent studies have examined the wage elasticity of job applications (e.g., Azar et al., 2022; Banfi and Villena-Roldan, 2019) and recruitment (e.g., Bassier et al., 2022; Hirsch et al., 2022). We add to this literature by showing how measures of posted pay premiums and amenities relate to vacancy duration.
@@ 2.1 Vacancy Data
We have access to almost 3 million job ads covering the near universe of publicly posted vacancies in Norway between 2002 and 2019. These data are maintained by the Norwegian Public Employment Agency (NAV). The employment agency collects information about vacancies from several sources, including online job boards and newspapers, as well as vacancies that are directly reported by employers to the agency.101010 In accordance with the Labor Market Act §7, Norwegian employers are required by law to report publicly posted job ads to the agency, which maintains a comprehensive database of publicly posted vacancy ads with the stated goal of providing job-seekers with current information on suitable job opportunities. More than 80 percent of job postings recorded in the agency’s database were retrieved from various online job boards. [^11] Bhuller et al. (2023) study the consequences on labor market matching of increased online job search and recruitment triggered by a roll-out of broadband internet across Norway during the early 2000s. The remaining 20 percent were either (i) scanned or transcribed from newspapers by caseworkers in the employment agency, or (ii) based on notifications of job openings sent by employers directly to the agency. The ability to observe virtually all publicly posted vacancy ads in the economy with full-text corpora is an important advantage of our setting, which differentiates it from the existing literature using vacancy texts that often relies on information from selected online job portals, typically covering job ads posted over a shorter time span.
@@
In addition to the job title and actual text of each job opening, our data contain the following structured information about each ad: unique establishment identifiers, the dates when the ad was registered and filled or removed (i.e., vacancy duration), the number of job openings per vacancy, and some additional information about job characteristics submitted by employers. The establishment identifiers are central to our analysis. They allow us to link each job opening to matched employer-employee administrative data and to compare the information we extract from the text to actual outcomes at the establishment level.121212 Notably, around 7 percent of job ads in our data were posted through recruitment or temporary employment agencies, and we drop these from most of our analysis, as we are unable to link such ads to the actual establishment where the job is placed. Further, in around 13 percent of job ads, the posting employer has for various reasons decided not to disclose the establishment name in the publicly posted information, but we do have the corresponding establishment identifier in our data as the agency maintains this information and could share this with us for research purposes (by the Labor Market Act §7-4). Each job ad in our data set also has an occupational code based on the 4-digit ISCO classification. This occupational code is assigned by caseworkers based on the job title and textual information on job descriptions and skill requirements stated by the employers in the vacancy, and is a novel feature of our data. Besides the structured information, the text of a job ad typically contains about 200 words. As we describe in Section 3, we use tools from the natural language processing literature to retrieve information on job attributes from these texts.
@@ 2.2 Matched Employer-Employee Data
We also have access to Norwegian administrative matched employer-employee data for all years between 2000 and 2019. The data come as a series of employment spells in each year, where each spell has information on the individual’s employer, as well as their pre-tax earnings, hours of work, and employment start and end date. Individuals can have several recorded spells in any given month if they receive earnings from more than one employer. For each worker, we aggregate all employment spells to the annual level and retain observations corresponding to the main employer, defined as the establishment with the largest annual earnings. Using information on annual earnings and annual contracted hours of work, we further calculate the average hourly wage for each worker in their main job in each year.
This data set has additional information on several background characteristics of employees (gender, education) and employers (location, industry). In the most recent years, the data set also has some information on a small subset of job attributes (e.g., shift work, contract type) for each employment spell that employers must report to Statistics Norway.
@@ 2.3 Sample Selection and Analysis Periods
While our vacancy and employer-employee data sets cover almost two decades, both in our descriptive and structural analyses we split the data into four five-year periods: 2000-2004, 2005-2009, 2010-2014, and 2015-2019. We do this in order to capture potential changes in workers’ valuation of employers in terms of their pay and non-pay attributes, as well as employers’ posting behavior. In most of our analysis, we focus on the 2015-2019 period as our baseline and provide supplementary results for the remaining five-year periods.
We further impose several restrictions that limit our focus to establishments for which we observe both job ads and job-to-job flows. Table 1 describes these restrictions and shows how they reduce the number of observations used in our baseline analysis for the 2015-2019 period. [^13] Appendix Table A.1 documents these sample restrictions for each of the four periods used in the analysis, as well as the 2000-2019 period as a whole. Our initial matched employer-employee data contains about 300,000 establishments covering about 3.1 million unique workers and 13.3 million worker-year observations. During the 2015-2019 period, we observe about 900,000 job ads. [^14] An ad is on average associated with 1.6 job openings, so these ads cover 1.4 million job openings. First, we limit our analysis to job ads with complete vacancy texts and to establishments that posted at least one job ad during the corresponding five-year period. [^15] As we explain in Section 4, we also group similar establishments using a k-means clustering approach. Appendix Table A.2 shows that restricting the sample strictly to establishments with observed job ad texts would only slightly lower our sample sizes. However, using the clustering approach greatly reduces the dimension of the number of employer-specific parameters we need to estimate in our structural analysis. For almost 60 percent of establishments observed in our employer-employee data, we do not observe any job ad texts in the corresponding five-year period. Despite this restriction, the remaining establishments cover almost 90 percent of worker-year observations, reflecting that posting establishments are larger. [^16] See Appendix Table A.3 for a comparison of establishment characteristics for different sample restrictions. Second, we remove job ads posted by staffing agencies, as for such ads we observe the establishment identifier for the staffing agency and not the actual employer. [^17] Around 92.5 percent of job ads classified as “staffing agency” ads in our data were posted by temporary employment agencies (e.g., Manpower), while recruitment agencies account for the remaining share. Third, our structural estimation is further restricted to establishments in the set of strongly connected establishments, i.e., the set of establishments across which we observe two-way job mobility flows during the corresponding five-year period. This restriction alone excludes almost 75 percent of establishments observed in our employer-employee data, but more than 80 percent of worker-year observations are still retained. Lastly, when we relate the employer-level model estimates to the information extracted from job ads, we impose both restrictions simultaneously. Interestingly, in this sample, we still observe job ads for nearly all establishments in the strongly connected set.
@@
All Observations $915,790$ (100%) $301,716$ (100%) $3,123,055$ (100%) $13,278,335$ (100%)
Observe Vacancy Text $863,478$ (94.3%) $127,913$ (42.4%) $2,733,050$ (87.5%) $11,898,501$ (89.6%)
Excluding Staffing Agencies $774,076$ (84.5%) $127,904$ (42.4%) $2,732,297$ (87.5%) $11,895,973$ (89.6%)
Strongly Connected Set $744,108$ (81.3%) $78,133$ (25.9%) $2,500,325$ (80.1%) $10,977,778$ (82.7%)
All Sample Restrictions $624,992$ (68.2%) $78,118$ (25.9%) $2,499,506$ (80.0%) $10,974,896$ (82.7%)
Note: This table documents the sample sizes for different sample restrictions imposed in the estimation for the 2015-2019 baseline period. “Observe Vacancy Text” is the set of vacancies with non-missing text content and at least one section written in Norwegian. “Excluding Staffing Agencies” is the subset of ads with observed text that are not posted by recruitment or temporary employment agency. “Strongly Connected Set” refers to the strongly connected set of employers used in the structural estimation. Sample sizes are calculated after linking job ads to matched employer-employee data at the establishment-cluster level. Appendix Tables A.1-A.2 provide sample selection for each of the four five-year periods (2000-2004, 2005-2009, 2010-2014, and 2015-2019) and the combined 2000-2019 period, with linking job ads to matched employer-employee data at the cluster and the establishment level, respectively.

### v2

@@ Abstract
First Version: July 2024
[…]
Keywords: amenities, non-pecuniary job attributes, compensating differentials, worker mobility, information frictions, pay transparency, job ads, text analysis, vacancy duration
JEL: J23, J32, J33, J62, J63
@@ 1 Introduction
There is a long tradition in labor economics of seeing jobs as more than pay (e.g., Brown, 1980; Duncan and Holmlund, 1983; Rosen, 1986). Phrases commonly used in the literature such as “compensating differentials,” “amenities,” or “workplace differentiation” all conceptualize the notion that jobs can be represented as a multidimensional bundle of attributes that workers trade off when considering opportunities with alternative employers. Such trade-offs can be substantial. Recent experimental evidence shows that workers at a US call center are willing to give up 8 percent of their salary for the option to work from home (Mas and Pallais, 2017), while undergraduates at a top US university would accept a 5.1 percent salary cut for a job offering the possibility to work part-time (Wiswall and Zafar, 2018).
What is unclear from this evidence, however, is how workers gather information on the pay and amenity attributes of a job while exploring their options in the labor market. The content of job postings represents one important search channel through which workers can gather information about these attributes. [^4] Recently, Carrillo-Tudela et al. (2023) provide survey evidence from Germany showing that 55.3 percent of firms used job postings in their latest hire, while 88.1 percent of workers used job postings in their search process. Notably, however, job postings are not exclusive of other search methods, such as social networks. Carrillo-Tudela et al. (2023) find that firms and workers use, respectively, 1.9 and 2.3 channels on average. But as employers ultimately decide on the content of publicly advertised job postings, it remains an open question whether such content in fact reflects the actual pay and amenities associated with different employers.
@@
To characterize the distribution of job attributes advertised by employers in job ads, we use the near-universe of publicly posted vacancies in Norway across several years. A key contribution of our paper is to systematically extract job attributes from these texts. Using tools from natural language processing, we identify nearly fifty commonly advertised pay and non-pay attributes, such as “competitive pay,” “flexible work hours,” and “nice work environment.” We build a large collection of unique natural language expressions associated with each attribute, which we make publicly available online for other researchers to use. [^5] The full list of expressions is available in an online repository, accessible through this link. This approach allows us to identify a large number of job attributes and to transparently map each attribute to many alternative expressions.
@@
A key strength of our data set is that each ad can be directly traced to the establishment posting the job. To shed light on how different employers advertise different types of information in publicly posted job ads and to assess the information content of these publicly advertised job attributes, we estimate workers’ valuation of employers using job flows and hourly wages derived from matched employer-employee data. Building on Sorkin (2018) and Morchio and Moser (2024), we use a structural revealed preference approach and obtain measures of workers’ valuation of different employers. Within our framework, this valuation can be further decomposed into a pay component, a non-pay component, and a job security component. Measuring these components separately matters because many of the attributes advertised in job listings, such as workplace quality, are not necessarily reflected in pay.
@@
We show that the content of job ads has predictive power for workers’ valuation of employers. Notably, we find that the attributes we detect in job ads can explain about 60 percent of the variation in overall employer values and employer pay premiums, and above 50 percent of amenity values and 40 percent of job security differences across employers. Job ads have predictive power over and above commonly observed job characteristics. Once we remove variation in employer values that can be attributed to flexible industry, occupation, and location controls, job attributes listed in ads can still explain 15-20 percent of the remaining variation across employers. In this regard, we find that attributes related to having flexible working hours and shift work are especially important predictors. Further, controlling for industry, occupation and location, we show that employers with higher pay premiums and amenity values associated with the information posted in job ads have shorter duration of posted vacancies, consistent with such employers being more attractive.
Finally, we use our model estimates to gauge the impact on worker mobility of having access to the content of job ads. We predict the counterfactual value of employers under alternative assumptions on workers’ information sets and derive the implied mobility flows in each case. In our baseline, we find that aggregate job-to-job transitions to more highly-valued employers increase by 1.1 percent in an economy where workers have access to information on job attributes in job ads, as compared to a benchmark economy where only “one-liner” vacancies listing occupation, industry, and location are posted. This increase corresponds to 25 percent of the additional mobility towards better employers in the full-information model. Our results further suggest that the information in job ads can contribute to men and women self-selecting into distinct employers based on their preferences.
@@
Our results also have implications for the design of policies related to the information contained in job ads. There is a current policy drive to improve pay transparency within organisations in many countries, and several jurisdictions are also implementing policies to promote pay transparency in job openings Cullen (2024). [^6] As an example, in November 2022 New York City mandated employers to systematically include a salary bracket in their job ads. Taken together, our results on the prevalence of specific job attributes in vacancy texts and on the type of attributes advertised by more attractive employers can be helpful to guide further policy efforts in that direction.
@@
We contribute to this literature by instead considering a broad set of pay and non-pay attributes that are advertised by employers in job listings. In a similar vein, several recent contributions focus on a single specific attribute, such as the type of flexible work arrangements advertised in job ads (Adams-Prassl et al., 2023a), the possibility to work remotely (Hansen et al., 2023), and whether the employer offers training on the job (Adams-Prassl et al., 2023b). We add to these important contributions by retrieving a comprehensive set of job attributes advertised in job ads, which encompasses flexible work arrangements, remote work and on-the-job training, but also what employers communicate about compensation or the quality of the workplace. In a similar spirit, Sockin (2022) uses text analysis to retrieve information on a large set of amenities from employer reviews posted by workers, rather than using vacancy texts. We see our results using vacancy texts as complementary since they instead reflect the information advertised by employers. [^7] In other recent and related contemporary work, Arold et al. (2024) and Lagos (2024) use text analysis methods to retrieve information on amenities contained in collective bargaining agreement texts.
@@
A large body of theoretical and empirical work has emphasized that non-pay attributes are important determinants of how much workers value alternative employers. One strand of this literature elicits workers’ preferences over specific attributes in survey experiments or quasi-experimental settings. Examples of non-pay attributes analyzed using this approach include workplace flexibility (Mas and Pallais, 2017; Wiswall and Zafar, 2018; Drake et al., 2023), job security (Datta, 2019), commuting time (Le Barbanchon et al., 2021), working from home (Nagler et al., 2022; Lewandowski et al., 2022), and shift work (Desiere and Walter, 2023), among others. Our paper instead provides a composite measure of employer amenity value that is linked to a large number of job attributes from job ads.
@@
We follow the second strand of this literature, which recovers structural estimates of the amenity value of employers by modelling where workers choose to work in a labor market with heterogeneous employers. In this approach, estimates of the amenity value of employers are obtained by reconciling workers’ choice of employer with their level of pay (Rosen, 1986). The choice of workers are captured using either the size of employers in a static setting (e.g., Card et al., 2018; Lamadon et al., 2022) or mobility flows between employers (e.g., Sorkin, 2018; Taber and Vejlin, 2020; Morchio and Moser, 2024). We rely on the latter approach since it can be recast in a job search framework, allowing us to further disentangle the amenity value of employers from differences in their recruitment effort. Relative to these important studies, we link the estimates of workers’ valuation of employers to the attributes advertised in job openings and quantify the value of this information. [^8] Morchio and Moser (2024) also relate their estimates of the amenity values of employers to some of the workplace characteristics observed in their data, such as “workplace hazards” and “working hours flexibility”.
@@
We also contribute to the expanding body of work that aims at unpacking search frictions in the labor market. Jäger et al. (2024) elicit the beliefs of workers on their reemployment wage if they were to lose their job and compare these wages to actual labor market outcomes. Horton et al. (2024) and Belot et al. (2022) design experiments within a controlled environment to study how job seekers respond to variation in the wage featured in job vacancies. Our approach adds to these studies in that we focus on the texts of actual job postings as a source of information on alternative employers. Our evidence also relates to recent studies on how posted wages relate to the duration of vacancies (e.g., Mueller et al., 2024; Bassier et al., 2023; Faberman and Menzio, 2018). [^9] Several recent studies have also examined the elasticity of applications (e.g., Azar et al., 2022; Banfi and Villena-Roldan, 2019) and hires (e.g., Bassier et al., 2022; Hirsch et al., 2022) with respect to posted wages. We add to this literature by showing how measures of posted pay premiums and amenities relate to the duration of vacancies.
@@ 2.1 Vacancy Data
We have access to almost 3 million job ads covering the near universe of publicly posted vacancies in Norway between 2002 and 2019. These data are maintained by the Norwegian Public Employment Agency (NAV). The employment agency collects information about vacancies from several sources, including online job boards and newspapers, as well as direct job opening reports from employers.101010 In accordance with the Labor Market Act §7, Norwegian employers are required by law to report publicly posted job openings to the agency, which maintains a comprehensive database of publicly posted vacancy ads with the stated goal of providing job-seekers with current information on suitable job opportunities. The share of online postings has increased gradually over the two decades, and about 80 percent of job postings recorded in the agency’s database since 2015 were retrieved from various online job boards. [^11] Bhuller et al. (2023) study the consequences on labor market matching of increased online job search and recruitment triggered by a roll-out of broadband internet across Norway during the early 2000s. The remaining job ads were either (i) scanned or transcribed from newspapers by caseworkers in the employment agency, or (ii) enclosed in the notifications of job openings sent by employers directly to the agency. The ability to observe virtually all publicly posted vacancy ads in the economy with full-text corpora is an important advantage of our setting, which differentiates it from the existing literature using vacancy texts that often relies on information from selected online job portals, typically covering job ads posted over a shorter time span.
@@
In addition to the job title and actual text of each job opening, our data contain the following structured information about each ad: unique establishment identifiers, the dates when the ad was registered and filled or removed (i.e., vacancy duration), the number of job openings per vacancy, and some additional information about job characteristics submitted by employers. The establishment identifiers are central to our analysis. They allow us to link each job opening to matched employer-employee administrative data and to compare the information we extract from the text to actual outcomes at the establishment level.121212 Notably, around 7 percent of job ads in our data were posted through recruitment or temporary employment agencies, and we drop these from most of our main analysis, as we are unable to link such ads to the actual establishment where the job is placed. Further, in around 13 percent of job ads, the posting employer has for various reasons decided not to disclose the establishment name in the publicly posted information, but we do have the corresponding establishment identifier in our data as the agency maintains this information and could share this with us for research purposes (according to the Labor Market Act §7-4). Each job ad in our data set also has an occupational code based on the 4-digit ISCO classification. This occupational code is assigned by caseworkers based on the job title and textual information on job descriptions and skill requirements stated by the employers in the vacancy, and is a novel feature of our data. Besides the structured information, the text of a typical job ad contains about 200 words. As we describe in Section 3, we use tools from the natural language processing literature to retrieve information on job attributes from these texts.
@@ 2.2 Matched Employer-Employee Data
We also have access to Norwegian administrative matched employer-employee data for all years between 2000 and 2019. The data come as a series of employment spells in each year, where each spell has information on the individual’s employer, as well as their pre-tax earnings, hours of work, and employment start and end date. Individuals can have several recorded spells in any given month if they receive earnings from more than one employer. For each worker, we aggregate all employment spells to the annual level and retain observations corresponding to the main employer, defined as the establishment with the largest annual earnings. Using information on annual earnings and annual contracted hours of work, we further calculate the average hourly wage for each worker in their main job in each year. This data set has additional information on several background characteristics of employees (gender, education) and employers (location, industry). In the most recent years, the data set also has some information on a small subset of job attributes (e.g., shift work, contract type) for each employment spell that employers must report to Statistics Norway.
@@ 2.3 Sample Selection and Analysis Periods
While our vacancy and employer-employee data sets cover almost two decades, both in our descriptive and structural analyses we split the data into four five-year periods: 2000-2004, 2005-2009, 2010-2014, and 2015-2019. We do this in order to capture potential changes in workers’ valuation of employers in terms of their pay and non-pay attributes, as well as employers’ posting behavior. In most of our analysis, we focus on the 2015-2019 period as our baseline and provide supplementary results for the remaining three five-year periods.
We impose several sample restrictions at various stages of our analysis. Table 1 describes these restrictions and shows how they reduce the number of observations used in our baseline analysis for the 2015-2019 period. [^13] Appendix Table A.1 shows the sample restrictions for each five-year period, and jointly for 2000-2019. As shown in the first row, prior to the selection restrictions, our initial matched employer-employee data contains about 300,000 establishments covering about 3.1 million unique workers and 13.3 million worker-year observations. During this period, we observe about 900,000 job ads in the employment agency’s database. [^14] An ad is on average associated with 1.6 job openings, so these ads cover 1.4 million job openings.
In Panel A of Table 1, we focus on the sample used for our text analysis in Section 3, where we limit attention to job ads with complete (non-missing) vacancy texts with at least one section written in Norwegian. As shown in Column (1), this restriction drops about 6 percent of job ads. Note, however, that less than half of the establishments in our initial sample publicly posted any job ads in the 2015-2019 period. Once we consider job ads with non-missing Norwegian texts, we are left with about 41 percent of establishments, as shown in Column (2). Meanwhile, the posting establishments tend to be substantially larger, and thus we still retain establishments covering about 90 percent of workers and worker-year observations in our text analysis, as shown in Columns (3)-(4). In Appendix Table A.2, we provide a comparison of how the various sample restrictions impact the sample composition.
Next, in Panel B, we consider the sample used for the structural model estimation in Section 4. For the structural estimation, we focus on prime-aged workers, i.e., aged 20–60, following standard practice in the literature. This restriction alone drops about 15 percent of workers and worker-year observations and 8 percent of establishments. Further, we restrict the estimation to establishments that are in the set of strongly connected establishments, i.e., the set of establishments across which we observe two-way job mobility flows during the 2015-2019 period. With both of these restrictions, we exclude more than 70 percent of the establishments in our initial employer-employee data, but can still retain establishments covering about 75 (70) percent of workers (worker-year observations). Comparing the average characteristics of establishments used in the text analysis and the model estimation in Appendix Table A.2, we find that the two samples are broadly similar in terms of industry and worker composition, while establishments used in the model estimation are larger.
@@
All Observations $915,789$ (100%) $292,327$ (100%) $3,123,052$ (100%) $13,278,337$ (100%)
Panel A: Text Analysis
   Observe Vacancy Text $860,467$ (94.0%) $118,741$ (40.6%) $2,882,624$ (92.3%) $11,893,847$ (89.6%)
Panel B: Model Estimation
   Prime-age Workers $880,892$ (96.2%) $268,565$ (91.9%) $2,671,777$ (85.6%) $11,253,721$ (84.8%)
   Strongly Connected Set $744,133$ (81.3%) $78,136$ (26.7%) $2,353,355$ (75.4%) $9,436,842$ (71.1%)
Panel C: Main Analysis
   Observe Vacancy Text $860,467$ (94.0%) $118,741$ (40.6%) $2,882,624$ (92.3%) $11,893,847$ (89.6%)
   Prime-age Workers $829,137$ (90.5%) $118,107$ (40.4%) $2,486,782$ (79.6%) $10,182,255$ (76.7%)
   Strongly Connected Set $701,117$ (76.6%) $78,123$ (26.7%) $2,353,237$ (75.4%) $9,436,263$ (71.1%)
   Excluding Staffing Agencies $623,341$ (68.1%) $77,164$ (26.4%) $2,324,318$ (74.4%) $9,239,281$ (69.6%)
Note: This table documents the sample sizes for different sample restrictions imposed in the estimation for the 2015-2019 baseline period. The first row shows the total number of observations in our data on job ads (Column 1) and matched employer-employee data (Columns 2-4), prior to the various sample selection steps. “Observe Vacancy Text” refers to the set of job ads with non-missing vacancy text and at least one section written in Norwegian. “Prime-age Workers” are the set of workers between 20 and 60 (inclusive).“Strongly Connected Set” refers to the strongly connected set of employers used in the structural estimation. “Excluding Staffing Agencies” is the set of employers that are not recruitment or temporary employment agencies. Panels A, B and C show the sample restrictions imposed on the text analysis, the structural model estimation, and the main analysis, respectively. Each sample restriction is imposed sequentially within each panel. Appendix Table A.1 provides an overview of sample selection for each of the four five-year periods (2000-2004, 2005-2009, 2010-2014, and 2015-2019) and the combined 2000-2019 period, respectively.
Lastly, in Panel C, we focus on the sample used in our main analysis in Section 5 where we relate the employer-level model estimates to the information extracted from job ads. For this analysis, we impose both sets of restrictions used in the earlier stages. In addition, we also drop job ads posted by staffing agencies, as for such ads we observe the establishment identifier for the staffing agency and not the actual employer. [^15] Around 92.5 percent of job ads classified as “staffing agency” ads in our data were posted by temporary employment agencies (e.g., Manpower), while recruitment agencies account for the remaining share. As shown in the final row of Table 1, for the main analysis, we retain about 68 percent of job ads, 26 percent of establishments, and 74 (70) percent of workers (worker-year observations) in our initial data.